use the power method to estimatedominant eigenvalue and the corresponding eigenvector for the matrix A =[1 2 1,0 2 5 ,1 0 -1.Start with the vector[1 1 1] and do two iterations.


1
Expert's answer
2021-10-05T17:06:24-0400

"A=\begin{pmatrix}
 1 & 2&1 \\
 0 & 2&5\\
1&0&-1
\end{pmatrix}"

"\begin{vmatrix}
 1-\lambda & 2&1 \\
 0 & 2-\lambda&5\\
1&0&-1-\lambda
\end{vmatrix}=(1-\lambda)(2-\lambda)(-1-\lambda)+10-2+\lambda="

"=(2-3\lambda+\lambda^2)(-1-\lambda)+\lambda+8=-\lambda^3+2\lambda^2+\lambda-2+\lambda+8="

"=-\lambda^3+2\lambda^2+2\lambda+6=0"

"\lambda_1=3.207"

"x_1=Ax_0=\begin{pmatrix}
 1 & 2&1 \\
 0 & 2&5\\
1&0&-1
\end{pmatrix}\begin{pmatrix}
 1 \\
 1\\
1
\end{pmatrix}=\begin{pmatrix}
 4 \\
 7 \\
0
\end{pmatrix}"


"x_1=x_1/||x_1||=\begin{pmatrix}
 0.4961 \\
 0.8682 \\
0
\end{pmatrix}"


"x_2=Ax_1=\begin{pmatrix}
 1 & 2&1 \\
 0 & 2&5\\
1&0&-1
\end{pmatrix}\begin{pmatrix}
 0.4961 \\
 0.8682 \\
0
\end{pmatrix}=\begin{pmatrix}
 2.2325 \\
 1.7364 \\
0.4961
\end{pmatrix}"


"x_2=x_2/||x_2||=\begin{pmatrix}
 0.7775 \\
 0.6047 \\
0.1728
\end{pmatrix}"


"Ax_2=\begin{pmatrix}
 1 & 2&1 \\
 0 & 2&5\\
1&0&-1
\end{pmatrix}\begin{pmatrix}
 0.7775 \\
 0.6047 \\
0.1728
\end{pmatrix}=\begin{pmatrix}
 2.1597 \\
 2.0734 \\
0.6047
\end{pmatrix}"


Approximation to the dominant eigenvalue "\lambda_1=3.207" :

"\lambda_1=\frac{(Ax_2)\cdot x_2}{x_2\cdot x_2}=\frac{3.0374}{1.0019}=3.032"


Find corresponding eigenvector:

"\begin{cases}
 (1-\lambda)x+2y+z=0\\
 (2-\lambda)y+5z=0\\
x- (1+\lambda)z=0
\end{cases}"


"\begin{cases}
 -2.032x+2y+z=0\\
 -1.032y+5z=0\\
x-4.032z=0
\end{cases}"


"10.16x-11.032y=0"

"2y-7.193z=0"


"x=1.086y"

"y=3.6z"

"v_1=\begin{pmatrix}
 1 \\
 0.9 \\
0.25
\end{pmatrix}"